On my homework I have a couple of problems that have to do with conic sections. I have already done one in class but my teacher didnt explain it. The problem is x^2+y^2-3x+8y=20. Im stuck on what to do next.
Answer by scott8148(6628)

the x and y are both positive and squared, with equal coefficients; that means a circle

the general equation is ___ (x - h)^2 + (y - k)^2 = r^ ___ a circle, centered at (h,k), with radius r

you have to complete the squares for x AND y

x^2 - 3x + ? + y^2 + 8y + ?? = 20

x^2 - 3x + (3/2)^2 + y^2 + 8y + (8/2)^2 = 20 + (3/2)^2 + (8/2)^2

(x - 3/2)^2 + (y + 4)^2 = 38.25